Math Question

An investment of $93000 was made by a business club. The investment was split into three parts and lasted one year. The first part of the investment earned 8% interest, the second 6%, and the third 9%. Total interest from the investments was $7170. The interest from the first investment was six times the interest from the second. Find the amounts of the three parts of the investment.
Answer #1

Answer :

A + B + C = 93000

0.08A + 0.06B + 0.09C = 7170

0.08A = 6 * (0.06B) => A = 4.5B


Sub in the last into the first two to get

5.5B + C = 93000 => C = 93000 - 5.5B

0.3B + 0.09C = 7170


Sub in C = 93000 - 5.5B into 0.3B + 0.09C = 7170 to get

0.3B + 0.09(93000 - 5.5B) = 7170

0.3B+8370-0.495B = 7170

-0.195B = -1200


B = 6153.85

Sub to get C = 59153.85

and A = 27692.31